OpenStudy (anonymous):

A boat has a speed of 5 mph in calm water. It takes the boat 3 hours to travel upstream but only 2 hours to travel the same distance downstream. Using the data in the table, what is the value of c, the speed of the current?

OpenStudy (cwrw238):

its speed downstream is 5 + c mph upstream its 5 - c now use the formula speed = distance / time to get value of c

OpenStudy (cwrw238):

hope this helps

OpenStudy (cwrw238):

do you need more help?

OpenStudy (anonymous):

thanks, it did help

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

i need hel p with another ?

OpenStudy (anonymous):

Paul can install a 300-square-foot hardwood floor in 18 hours. Matt can install the same floor in 22 hours. How long would it take Paul and Matt to install the floor working together?

OpenStudy (cwrw238):

Paul can work at the rate of 300/18 sq ft per hour Matt can work at the rate of 300/22 sq ft per hour

OpenStudy (cwrw238):

be back in a short while...

OpenStudy (cwrw238):

sorry I was so long away actually you don't need to know the above rates of work solve this equation for x 1/18 + 1/22 = 1/x x is the required time